# LiRA: A Multi-Agent Framework for Reliable and Readable Literature Review Generation

### v2

@@ 2. RQ2
    How well are the generated articles written in terms compared to existing literature writing methods when all are given the same set of references?
@@ 2 Related Work
Comparisons have been made between agentic LLM systems and human cognition, as the idea of breaking down a task into smaller steps, which is often done by these workflows, is commonly used to describe how humans solve more intricate problems (flower_cognitive_1981; correa_humans_2023). Using this concept, several works show promise in implementing agentic workflows in various fields, such as the medical sciences (tang_medagents_2024) and law (watson_law_2025). In some cases, this results in an improvement of more than 90% compared to a simple baseline (watson_law_2025).
@@ 3.1 Outline Drafter Agent
To manage the context size and maintain focus on the most relevant content, the outline is constructed from up to 50 references. We use existing heuristics (wang_autosurvey_2024) that recommend generating approximately 8 sections with around 4 subsections, but the agent can adapt this amount as needed. We also include default sections such as an introduction and conclusion to ensure consistency and completeness.
@@ 3.2 Subsection Writer Agent
Relevant references are retrieved using the FAISS index (douze_faiss_2025) based on section-level descriptions, capped at 25% of the total pool per subsection (min 3, max 150). This balances citation diversity with input tractability. Moreover, the reference titles, abstracts, or full texts can be retrieved depending on the availability. The agent outputs 1̃,000 words per subsection, enabling long-form synthesis while staying within model context limits. Also, the article title, abstract, and conclusion are written after the body is generated, mimicking human writing practices and avoiding the premature commitment of top-down generation.
@@ 4.1 Baselines
Introduced in an automated survey-writing challenge (tian_overview_2025), MASS is the only agentic framework with publicly available code. Its workflow differs fundamentally from LiRA: Instead of decomposing the writing process into specialized roles with iterative feedback, MASS follows a strictly sequential pipeline. The system first clusters references by topical similarity to construct an outline, then generates section contents and a title directly from these clusters, and finally appends a conclusion. Long reference lists are handled by passing them as attachments.
@@ 4.4 Results
Table 1: Results for similarity to the human-written reviews with the baseline settings. Note that if both numbers in a row are bolded, it means they returned the exact same value.
@@
Figure 2: SME evaluation results.
@@ Results
The results for this setting can be found in figure 4. Though the extent of bias mitigation itself cannot be measured properly with the current metrics, we can remark how using gemma3:4b has little impact on the scores overall across all metrics. This suggests that alternative model configurations may not significantly alter the article output quality, therefore indicating potentially that the pipeline can work well even when using different LLMs in the process.
@@
Table 4: The gemma3 results for the ScienceDirect dataset.

### v3

@@ 2. RQ2
    How well written are the generated articles compared to existing literature writing methods when all are given the same set of references?
@@ 2 Related Work
Comparisons have been made between agentic LLM systems and human cognition, due to how breaking down a task into smaller steps, which these workflows often do, is commonly used to describe how humans solve more intricate problems (flower_cognitive_1981; correa_humans_2023). Using this concept, several works show promise in implementing agentic workflows in various fields, such as the medical sciences (tang_medagents_2024) and law (watson_law_2025). In some cases, this results in an improvement of more than 90% compared to a simple baseline (watson_law_2025).
@@ 3.1 Outline Drafter Agent
To manage the context size and focus on only the most relevant content, the outline is constructed from up to 50 references. We use existing heuristics (wang_autosurvey_2024) that recommend generating approximately 8 sections with around 4 subsections, but the agent can adapt this amount as needed. We also include default sections such as an introduction and conclusion to ensure consistency and completeness.
@@ 3.2 Subsection Writer Agent
Relevant references are retrieved using the FAISS index (douze_faiss_2025) based on section-level descriptions, capped at 25% of the total pool per subsection (min 3, max 150). This balances citation diversity with input tractability. Moreover, the reference titles, abstracts, or full texts can be retrieved depending on the availability. The agent outputs $\sim$1,000 words per subsection, enabling long-form synthesis while staying within model context limits. Also, the article title, abstract, and conclusion are written after the body is generated, mimicking human writing practices and avoiding the premature commitment of top-down generation.
@@ 4.1 Baselines
Introduced in an automated survey-writing challenge (tian_overview_2025), MASS is the only agentic framework from that challenge with publicly available code. Its workflow differs fundamentally from LiRA: Instead of decomposing the writing process into specialized roles with iterative feedback, MASS follows a strictly sequential pipeline. The system first clusters references by topical similarity to construct an outline, then generates section contents and a title directly from these clusters, and finally appends a conclusion. Long reference lists are handled by passing them as attachments.
@@ 4.4 Results
Table 1: Results for similarity to the human-written reviews with the baseline settings.
@@
Figure 2: SME evaluation results. Here, C indicates Coverage, S indicates Structure, and R indicates Relevance.
@@ Results
The results for this setting can be found in Table 4. Though the extent of bias mitigation itself cannot be measured properly with the current metrics, we can remark how using gemma3:4b has little impact on the scores overall across all metrics. This suggests that alternative model configurations may not significantly alter the article output quality, therefore indicating potentially that the pipeline can work well even when using different LLMs in the process.
@@
Table 4: The gemma3 results for the ScienceDirect dataset. Note that if both numbers in a row are bolded, it means they returned the exact same value.